Grade 12th passAlgebra

Find the whole quantity if
5% of it is 600.........

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let us assume the whole quantity be x,
Then,
(12/100) × (x) = 1080
X = 1080 × (100/12)
X = 540 × (100/6)
X = 90 × 100
X =₹9000

Thanks